OpenStudy (anonymous):

The femur bone in a human leg has a minimum effective cross section of 3.15cm^2. How much compressive force can it withstand before breaking? Assume the ultimate strength of the bone to be 1.70*10^8 N/m^2. Answer in N.

OpenStudy (anonymous):

Same as before, you are just given new information. Notice the units of the cross-section. This is an area A. Notice the units of the strength. This is a stress strength S. You are looking for force. S = F / A so: F = S * A